# Assimilative Causal Inference

### v1

@@ B.3 Calculation of the Subjective and Objective CIRs for CGNSs
Using the online smoother for CGNSs, for each $j\in\llbracket N\rrbracket$ we can compute both the subjective and objective CIRs over $[t_{j},T]$ for the potential assimilative causal link $\mathbf{y}^{j}\rightarrow\mathbf{x}$ (with analogous results for the CCIRs). Let $p_{n}(\mathbf{y}^{j}|\mathbf{x})$, for $j\in\llbracket N\rrbracket$ and $n\in\{j,\ldots,N\}$ denote the $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{s}}^{j,n},\mathbf{R}_{\text{s}}^{j,n})$ PDF from Theorem B.3. At each $t_{j}=j\Delta t$, we quantify the information gain in the complete smoother distribution $p_{N}(\mathbf{y}^{j}|\mathbf{x})$ beyond its lagged counterpart $p_{n}(\mathbf{y}^{j}|\mathbf{x})$ using the signal-dispersion decomposition from (1.6) as:
$\mathcal{P}^{j}_{n}:=\mathcal{P}\big{(}p_{N}(\mathbf{y}^{j}|\mathbf{x}),p_{n}(\mathbf{y}^{j}|\mathbf{x})\big{)}$ $=\frac{1}{2}\big{(}\boldsymbol{\mu}_{\text{s}}^{j,N}-\boldsymbol{\mu}_{\text{s}}^{j,n}\big{)}^{\mathtt{T}}(\mathbf{R}_{\text{s}}^{j,n})^{-1}\big{(}\boldsymbol{\mu}_{\text{s}}^{j,N}-\boldsymbol{\mu}_{\text{s}}^{j,n}\big{)}$ (2.8)
@@
Per Section AA.5 and (1.16), we identify $\delta(t_{n};t_{j})=\mathcal{P}^{j}_{n}$, for $j\in\llbracket N\rrbracket$ and $n\in\{j,\ldots,N\}$. This yields:
@@
    $\tilde{\tau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)=\max_{n\in\{j,\ldots,N\}}\big{\{}\mathcal{P}^{j}_{n}>\varepsilon\big{\}}\Delta t-t_{j}\in[0,T-t_{j}].$ (2.9)
@@
    $\tau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=\frac{\Delta t}{\max_{n}\{\mathcal{P}^{j}_{n}\}}\sum_{n=j}^{N}\mathcal{P}^{j}_{n}\in[0,T-t_{j}].$ (2.10)
We maintain the conventions from Section AA.5: $\tilde{\tau}=0$ when no maximum exists, and $\tau=\tau^{\text{approx}}=0=\tilde{\tau}$ when $\max_{n}\{\mathcal{P}^{j}_{n}\}=0$. For strongly intermittent systems and the objective CIR, we employ the computationally efficient scheme in (2.10) rather than the full integral definition.
@@
[…]
@@ C.1 Principle of Nil Assimilative Causality for CGNSs
[…]
@@
$p_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})=p_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x}),\quad t\in[0,T].$
[…]
@@
$\overleftarrow{\mathrm{d}(\boldsymbol{\mu}_{\text{{s}}}-\boldsymbol{\mu}_{\text{{f}}})}(t)=\overleftarrow{\mathrm{d}\boldsymbol{\mu}_{\text{{s}}}}(t)+\mathrm{d}\boldsymbol{\mu}_{\text{{f}}}(t)=-(\boldsymbol{\Lambda}^{\mathbf{y}}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})\mathbf{R}_{\text{{f}}}^{-1})(\boldsymbol{\mu}_{\text{{s}}}-\boldsymbol{\mu}_{\text{{f}}})\mathrm{d}t,\quad T\geq t\geq 0,$ (3.1)
@@
$\boldsymbol{\mu}_{\text{{s}}}(t)=\boldsymbol{\mu}_{\text{{f}}}(t),\quad t\in[0,T].$
@@
where in the last equality we have used the stability property of conditional expectations, $\mathcal{F}_{t}^{\mathbf{x}}\subseteq\mathcal{F}_{T}^{\mathbf{x}}$, and the $\mathcal{F}_{t}^{\mathbf{x}}$-measurability of $\mathbf{R}_{\text{{f}}}(t)$, with $\mathcal{F}_{t}^{\mathbf{x}}$ denoting the $\sigma$-algebra generated by $\{\mathbf{x}(s)\}_{s\leq t}$ for $t\in[0,T]$ [15]. Combining these, we end up with:
$p_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})=p_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x}),\quad t\in[0,T],$
[…]
@@ C.2 Principle of Nil Conditional Assimilative Causality for CGNSs
The following theorem extends Theorem C.1 to establish the principle of nil conditional assimilative causality for CGNSs. To formulate this result, we first introduce some necessary notation. Following Section AA.4, we consider the state variables $(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}},\mathbf{y})$ and reformulate the CGNS in (1.1) as:
@@
We also define the Schur complement of the $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ blocks with respect to $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ [7]:
@@
under the tacit assumption that $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ are invertible; $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ is positive definite if and only if either one of $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ or $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ and its associated Schur complement are positive definite matrices [7]. In the case where both $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ are positive definite, then $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}$ enjoys the following explicit representation:
@@
×$\hskip 56.9055pt\times\begin{pmatrix}\mathbf{I}_{k_{\text{{A}}}\times k_{\text{{A}}}}&-(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})^{-1}\\ -(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})^{-1}&\mathbf{I}_{k_{\text{{B}}}\times k_{\text{{B}}}}\end{pmatrix}.$
[…]
@@
[…]
@@
$p_{t}^{\text{{s}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}})=p_{t}^{\text{{f}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),\quad t\in[0,T].$
[…]
@@
For CGNS, the filter and smoother means evolve according to linear random ODEs, while their covariance matrices satisfy forward Riccati and backward symmetric Sylvester random ODEs respectively [1]. Under the regularity conditions of Theorems B.1 and B.2, these Gaussian statistics exhibit continuous dependence on both the model parameters in (2.1) and their initial/terminal conditions [8, 18]. This continuity allows direct computation of the posterior PDFs in (1.13) through the limit $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ applied to (2.2) and (2.4). Crucially, Gaussianity is preserved under this limit since a normal distribution depends continuously on its mean and covariance. Specifically, $p^{\text{f}|\mathbf{x}_{\text{B}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{A}})$ remains Gaussian $\mathcal{N}_{l}\big{(}\boldsymbol{\mu}_{\text{f}|\mathbf{x}_{\text{B}}}(t),\mathbf{R}_{\text{f}|\mathbf{x}_{\text{B}}}(t)\big{)}$, where:
@@
As the measurability of the posterior Gaussian statistics remains unaffected under the limit $\mathrm{Var}(\mathbf{x}_{\text{{B}}}(t))\to+\infty$, we just need to prove
$\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}(t)=\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}(t),\quad t\in[0,T],$
@@
×$\hskip 28.45274pt\times\begin{pmatrix}\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})/(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})\big{)}^{-1}&\mathbf{0}_{k_{\text{{A}}}\times k_{\text{{B}}}}\\ \mathbf{0}_{k_{\text{{B}}}\times k_{\text{{A}}}}&\big{(}(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})/(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})\big{)}^{-1}\end{pmatrix}$

### v2

@@ B.3 Calculation of the Subjective and Objective CIRs for CGNSs
Using the online smoother for CGNSs, for each $j\in\llbracket N\rrbracket$ we can compute both the subjective and objective CIRs over $t_{j},T]$ for the potential assimilative causal link $\mathbf{y}^{j}\rightarrow\mathbf{x}$ (with analogous results for the CCIRs). Let $p_{n}(\mathbf{y}^{j}|\mathbf{x})$, for $j\in\llbracket N\rrbracket$ and $n\in\{j,\ldots,N\}$ denote the $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{s}}^{j,n},\mathbf{R}_{\text{s}}^{j,n})$ PDF from Theorem B.3. At each $t_{j}=j\Delta t$, we quantify the information gain in the complete smoother distribution $p_{N}(\mathbf{y}^{j}|\mathbf{x})$ beyond its lagged counterpart $p_{n}(\mathbf{y}^{j}|\mathbf{x})$ using the signal-dispersion decomposition from (1.6) as:
$\mathcal{P}^{j}_{n}=\mathcal{P}\big(p_{N}(\mathbf{y}^{j}|\mathbf{x}),p_{n}(\mathbf{y}^{j}|\mathbf{x})\big)$ $=\frac{1}{2}\big(\boldsymbol{\mu}_{\text{s}}^{j,N}-\boldsymbol{\mu}_{\text{s}}^{j,n}\big)^{\mathtt{T}}(\mathbf{R}_{\text{s}}^{j,n})^{-1}\big(\boldsymbol{\mu}_{\text{s}}^{j,N}-\boldsymbol{\mu}_{\text{s}}^{j,n}\big)$ (2.8)
@@
Per Section A.5 and (1.16), we identify $\delta(t_{n};t_{j})=\mathcal{P}^{j}_{n}$, for $j\in\llbracket N\rrbracket$ and $n\in\{j,\ldots,N\}$. This yields:
@@
    $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)=\max_{n\in\{j,\ldots,N\}}\big\{\mathcal{P}^{j}_{n}>\varepsilon\big\}\Delta t-t_{j}\in[0,T-t_{j}].$ (2.9)
@@
    $\uptau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=\frac{\Delta t}{\max_{n}\{\mathcal{P}^{j}_{n}\}}\sum_{n=j}^{N}\mathcal{P}^{j}_{n}\in[0,T-t_{j}].$ (2.10)
We maintain the conventions from Section A.5: $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)=0$ when no maximum exists, and $\uptau_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=\uptau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=0=\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)$ when $\max_{n}\{\mathcal{P}^{j}_{n}\}=0$. For strongly intermittent systems and the objective CIR, we employ the computationally efficient scheme in (2.10) rather than the full integral definition.
In Algorithm 3 we provide a brief outline of how to calculate the subjective and objective CIR lengths of $\mathbf{y}^{j}\rightarrow\mathbf{x}$ over $t_{j},T]$ for $j\in\llbracket N\rrbracket$ in pseudocode, specifically for CGNSs. An analogous algorithm extends to CCIRs for CGNSs, as well as to the general complex system in (1.1) by appropriately computing $\delta(t_{n};t_{j})=\mathcal{P}^{j}_{n}$ for each $j\in\llbracket N\rrbracket$ and $n\in\{j,\ldots,N\}$.
Data: $N\in\mathbb{N}$, $0<\Delta t\ll 1$, $\{t_{j}=j\Delta t\}_{j\in\llbracket N\rrbracket}$, Observations of $\mathbf{x}$: $\{\mathbf{x}^{0},\ldots,\mathbf{x}^{N}\}$, $\boldsymbol{\mu}_{\text{{f}}}^{0}$, $\mathbf{R}_{\text{{f}}}^{0}$, $\varepsilon\geq 0$
Result: $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)$ for arbitrary $\varepsilon$, $\uptau_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}$, and $\uptau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}$, all for $j\in\llbracket N\rrbracket$
1 for $j\in\llbracket N\rrbracket$ do
2    for $n\in\{j,\ldots,N\}$ do
3        Calculate $p_{n}(\mathbf{y}^{j}|\mathbf{x})$, the Gaussian PDF corresponding to $\mathcal{N}_{l}(\boldsymbol{\mu}_{\text{s}}^{j,n},\mathbf{R}_{\text{s}}^{j,n})$, per Theorem B.3;
4
5    end for
6   Calculate $\delta(t_{n};t_{j})=\mathcal{P}^{j}_{n}$ using (2.8) for each $n\in\{j,\ldots,N\}$;
7    if $\varepsilon<\max_{n}\{\mathcal{P}^{j}_{n}\}$ then
8       $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)=\max_{n\in\{j,\ldots,N\}}\big\{\mathcal{P}^{j}_{n}>\varepsilon\big\}\Delta t-t_{j}$;
9   else
10       $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)=0$;
11    end if
12
13   if $\max_{n}\{\mathcal{P}^{j}_{n}\}>0$ then
14       if $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)$ is known for many $\varepsilon\in\big[0,\max_{n}\{\mathcal{P}^{j}_{n}\}\big]$ then
15           $\uptau_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=\frac{1}{\max_{n}\{\mathcal{P}^{j}_{n}\}}\int_{0}^{\max_{n}\{\mathcal{P}^{j}_{n}\}}\overset{\sim}{\uptau}_{\mathbf{y}(t)\rightarrow\mathbf{x}}(\varepsilon){\mathrm{d}}\varepsilon$;
16        end if
       $\uptau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=\frac{\Delta t}{\max_{n}\{\mathcal{P}^{j}_{n}\}}\sum_{n=j}^{N}\mathcal{P}^{j}_{n}$;
        /*/ If $\max_{n}\{\mathcal{P}^{j}_{n}\}$ is essentially negligible, e.g., $\max_{n}\{\mathcal{P}^{j}_{n}\}<10^{-4}$, we can set $\uptau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=0$, which necessarily coincides with $\uptau_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}\approx 0$, as to avoid operationally inflated values due to numerical imprecision
17
18   else
19       $\uptau_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=\uptau^{\text{approx}}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}=0$;
20    end if
21
22 end for
Algorithm 3 CIRs of $\mathbf{y}^{j}\rightarrow\mathbf{x}$ over $t_{j},T]$ for $j\in\llbracket N\rrbracket$ in the case of a CGNS
@@ 10       $\overset{\sim}{\uptau}_{\mathbf{y}(t_{j})\rightarrow\mathbf{x}}(\varepsilon)=0$;
Before investigating complex turbulent systems, we first validate the ACI framework’s theoretical foundations using the analytically tractable CGNS structure. Crucially, any reliable causal metric must satisfy two requirements: (1) It should be asymmetric in its state components, sometimes referred to as Schreiber’s principle [schreiber2000measuring], and (2) it should correctly identify unidirectional relationships, yielding zero measure in non-causal directions; in other words, it must obey the principle of nil causality [liang2016information], which is “an event cannot cause another if their dynamics are independent”. Both ACI and its conditional variant naturally satisfy (1) for general complex systems. As for the principle of nil causality, for CGNSs this principle takes a rigorous form: when $\mathbf{x}$’s evolution is independent of $\mathbf{y}$, the ACI framework guarantees $\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}$ for all $t\in[0,T]$, confirming the absence of spurious causality. This extends to conditional ACI, mutatis mutandis.
@@ C.1 Principle of Nil Assimilative Causality for CGNSs
$\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x},\qquad\forall\ t\in[0,T],$
@@
$p_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})=p_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x}),\qquad t\in[0,T].$
Note on Theorem C.1: The condition $\boldsymbol{\Lambda}^{\mathbf{x}}\equiv\mathbf{0}_{k\times l}$ prohibits $\mathbf{y}$ from entering the mean dynamics of $\mathbf{x}$, while $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})\equiv\mathbf{0}_{l\times k}$ nullifies the cross-interactions between their noise feedbacks. The combination of these two assumptions removes the possibility of $\mathbf{y}$ contributing to the evolution of $\mathbf{x}$ in the dynamics, both through its drift coefficient, as well as via the full diffusion coefficient of the system in (2.1). As a result: $\mathbf{y}\ \cancel{\rightarrow}\ \mathbf{x}$.
@@
$\overleftarrow{\mathrm{d}(\boldsymbol{\mu}_{\text{{s}}}-\boldsymbol{\mu}_{\text{{f}}})}(t)=\overleftarrow{\mathrm{d}\boldsymbol{\mu}_{\text{{s}}}}(t)+\mathrm{d}\boldsymbol{\mu}_{\text{{f}}}(t)=-(\boldsymbol{\Lambda}^{\mathbf{y}}+(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{y}})\mathbf{R}_{\text{{f}}}^{-1})(\boldsymbol{\mu}_{\text{{s}}}-\boldsymbol{\mu}_{\text{{f}}})\mathrm{d}t,\qquad T\geq t\geq 0,$ (3.1)
@@
$\boldsymbol{\mu}_{\text{{s}}}(t)=\boldsymbol{\mu}_{\text{{f}}}(t),\qquad t\in[0,T].$
@@
where in the last equality we have used the stability property of conditional expectations, $\mathcal{F}_{t}^{\mathbf{x}}\subseteq\mathcal{F}_{T}^{\mathbf{x}}$, and the $\mathcal{F}_{t}^{\mathbf{x}}$-measurability of $\mathbf{R}_{\text{{f}}}(t)$, with $\mathcal{F}_{t}^{\mathbf{x}}$ denoting the $\sigma$-algebra generated by $\{\mathbf{x}(s)\}_{s\leq t}$ for $t\in[0,T]$ [liptser2001statistics]. Combining these, we end up with:
$p_{t}^{\text{{s}}}(\mathbf{y}|\mathbf{x})=p_{t}^{\text{{f}}}(\mathbf{y}|\mathbf{x}),\qquad t\in[0,T],$
which by (1.4) yields $\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}$ for each $t\in[0,T]$. ∎
@@ C.2 Principle of Nil Conditional Assimilative Causality for CGNSs
The following theorem extends Theorem C.1 to establish the principle of nil conditional assimilative causality for CGNSs. To formulate this result, we first introduce some necessary notation. Following Section A.4, we consider the state variables $(\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}},\mathbf{y})$ and reformulate the CGNS in (1.1) as:
@@
We also define the Schur complement of the $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ blocks with respect to $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ [gallier2019notes]:
@@
under the tacit assumption that $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ are invertible; $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ is positive definite if and only if either one of $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ or $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ and its associated Schur complement are positive definite matrices [gallier2019notes]. In the case where both $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})$ and $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})$ are positive definite, then $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})^{-1}$ enjoys the following explicit representation:
@@
$\hskip 56.9055pt\times\begin{pmatrix}\mathbf{I}_{k_{\text{{A}}}\times k_{\text{{A}}}}&-(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})^{-1}\\ -(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})^{-1}&\mathbf{I}_{k_{\text{{B}}}\times k_{\text{{B}}}}\end{pmatrix}$.
With these preliminaries, we now establish the principle of nil conditional assimilative causality for CGNSs: when (i) $\mathbf{x}_{\text{A}}$’s evolution is independent of $\mathbf{y}$, and (ii) the uncertainty levels of $\mathbf{x}_{\text{A}}$ and $\mathbf{x}_{\text{B}}$ are non-interacting, the generalized ACI framework correctly identifies the absence of conditional causation. Formally, $\big(\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}_{\text{{A}}}\big)\big|\mathbf{x}_{\text{{B}}}$, for all $t\in[0,T]$.
@@
$\big(\mathbf{y}(t)\ \cancel{\rightarrow}\ \mathbf{x}_{\text{{A}}}\big)\big|\mathbf{x}_{\text{{B}}},\qquad\forall\ t\in[0,T],$
@@
$p_{t}^{\text{{s}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}})=p_{t}^{\text{{f}}|\mathbf{x}_{\text{{B}}}}(\mathbf{y}|\mathbf{x}_{\text{{A}}}),\qquad t\in[0,T].$
Note on Theorem C.2: The assumption $\boldsymbol{\Lambda}^{\mathbf{x}_{\text{A}}}\equiv\mathbf{0}_{k_{\text{A}}\times l}$ eliminates direct influence of $\mathbf{y}$ on $\mathbf{x}_{\text{A}}$’s mean dynamics, while $(\boldsymbol{\Sigma}^{\mathbf{y}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{A}}})\equiv\mathbf{0}_{l\times k_{\text{B}}}$ removes their noise feedback coupling. However, these conditions alone are insufficient to prevent indirect $\mathbf{y}$-$\mathbf{x}_{\text{A}}$ interactions through $\mathbf{x}_{\text{B}}$ via observable noise cross-correlations. We therefore additionally impose $(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{A}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{B}}})\equiv\mathbf{0}_{k_{\text{A}}\times k_{\text{B}}}$, rendering $(\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})$ block-diagonal. Collectively, these ensure: $\big(\mathbf{y}\ \cancel{\longrightarrow}\ \mathbf{x}_{\text{{A}}}\big)\big|\mathbf{x}_{\text{{B}}}$.
@@
For CGNS, the filter and smoother means evolve according to linear random ODEs, while their covariance matrices satisfy forward Riccati and backward symmetric Sylvester random ODEs respectively [kandil2003matrix]. Under the regularity conditions of Theorems B.1 and B.2, these Gaussian statistics exhibit continuous dependence on both the model parameters in (2.1) and their initial/terminal conditions [han2017random, neckel2013random]. This continuity allows direct computation of the posterior PDFs in (1.13) through the limit $\mathrm{Var}(\mathbf{x}_{\text{B}}(t))\to+\infty$ applied to (2.2) and (2.4). Crucially, Gaussianity is preserved under this limit since a normal distribution depends continuously on its mean and covariance. Specifically, $p^{\text{f}|\mathbf{x}_{\text{B}}}_{t}(\mathbf{y}|\mathbf{x}_{\text{A}})$ remains Gaussian $\mathcal{N}_{l}\big(\boldsymbol{\mu}_{\text{f}|\mathbf{x}_{\text{B}}}(t),\mathbf{R}_{\text{f}|\mathbf{x}_{\text{B}}}(t)\big)$, where:
@@
As the measurability of the posterior Gaussian statistics remains unaffected under the formal limit $\mathrm{Var}(\mathbf{x}_{\text{{B}}}(t))\to+\infty$, we just need to prove
$\boldsymbol{\mu}_{\text{{s}}|\mathbf{x}_{\text{{B}}}}(t)=\boldsymbol{\mu}_{\text{{f}}|\mathbf{x}_{\text{{B}}}}(t),\qquad t\in[0,T],$
@@
$\hskip 28.45274pt\times\begin{pmatrix}\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})/(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{B}}}})\big)^{-1}&\mathbf{0}_{k_{\text{{A}}}\times k_{\text{{B}}}}\\ \mathbf{0}_{k_{\text{{B}}}\times k_{\text{{A}}}}&\big((\boldsymbol{\Sigma}^{\mathbf{x}}\circ\boldsymbol{\Sigma}^{\mathbf{x}})/(\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}}\circ\boldsymbol{\Sigma}^{\mathbf{x}_{\text{{A}}}})\big)^{-1}\end{pmatrix}$